OpenStudy (anonymous):

giving medal for answer Find the number of sides in the polygon with whose angles measures have the given sum. 900° a. 4 c. 7 b. 5 d. 9

OpenStudy (ivettef365):

the formula to find the measure of the angles is (n-2)180

OpenStudy (ivettef365):

so you have 900, to go backwards on the formula first divide 900/180 then add 2

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

The formula for the sum of the measures of the interior angles of a regular polygon is (n-2)180. Set this equal to 900 and solve for n to find the number of sides in the polygon. 900 = 180(n-2) 900 = 180n - 360 180n = 1260 n = 1260/180 n = 7

OpenStudy (ivettef365):

yw :)